OpenStudy (anonymous):

simplify the expression: -(7)to the power of -2

OpenStudy (anonymous):

-7 ^-2 = 1/-7^2 = 1/49 = 0.02041...

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

negative powers take a number to a fraction, and the fraction is the same number with a positive power, if that makes any sense. i was going to draw it out for you but it's not working on my computer. you get it, right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

okay good

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can u help me with this one

OpenStudy (anonymous):

it depends, i might. whats the problem?

OpenStudy (anonymous):

3m^8p^0

OpenStudy (anonymous):

if you drew that out would it be 3m to the 8p power to the 0 power?

OpenStudy (anonymous):